OpenStudy (anonymous):

|x+5|=10 absoulute value

OpenStudy (jdoe0001):

\(\bf |x+5|=10\implies \begin{cases} +(x+5=10)\\ \quad \\ \bf -(x+5=10) \end{cases}\)

OpenStudy (ranga):

If |x| = a, then the solution is x = -a or x = a Can you use the above formula to solve |x + 5| = 10?

OpenStudy (anonymous):

so will x= 15

OpenStudy (ranga):

First solve x + 5 = 10 and then solve x + 5 = -10 Solve x + 5 = 10. What is x?